# Forest-Chat: Adapting Vision-Language Agents for Interactive Forest Change Analysis

### v1

@@ 2.3 Vision-Language Models (VLMs)
Vision–language models (VLMs) have seen rapid progress alongside recent advances in LLMs such as GPT (openai et al., 2023) and LLaMA (Touvron et al., 2023), combining instruction-following with visual feature analysis. VLMs typically pair a pretrained visual encoder — such as the contrastively trained CLIP model (Radford et al., 2021), or ViTs — with an LLM through a projection or transformation module that maps visual embeddings into the language model’s input space (Li et al., 2024d; Tao et al., 2025). This alignment allows LLMs to link user instructions to extract visual information, providing support for generalised tasks such as visual-question-answering, captioning, and zero-shot transfer learning (Lu et al., 2025; Li et al., 2024b). These capabilities have made VLMs an increasingly common choice for RS applications, particularly through large foundation models pretrained on massive and diverse datasets using task-agnostic objectives (Li et al., 2024c; Zhang et al., 2025b; Liu et al., 2024; Irvin et al., 2025; Liu and Lian, 2024; Li et al., 2024d; Hu et al., 2025; Zhang et al., 2024b). RS applications benefit from domain-specific pretraining across varying sensors, resolutions, and imaging geometries, while self-supervised and zero-shot learning helps reduce the reliance on exhaustive annotation (Peng et al., 2025; Lu et al., 2025). Text prompting and question-and-answer interactions further support context-aware interpretation, helping models link ambiguous visual patterns to high-level semantic concepts (e.g. deforestation) (Peng et al., 2025; Li et al., 2024a).
@@
Despite strong performance on natural imagery, general-purpose multimodal LLMs such as GPT-4V and LLaVa (openai et al., 2023; Bazi et al., 2024) perform poorly on RS and change-related tasks due to domain mismatch Zhang and Wang (2024). One response has been to adapt LLM-driven, task-specific spatio-temporal VLMs to RS vision-language understanding tasks (Liu et al., 2025). Examples include KCFI (Yang et al., 2025a) and Semantic-CC (Zhu et al., 2024), which fine-tune visual modules or lightweight LLM components for change-focused tasks. Semantic-CC additionally integrates SAM-derived change masks (Kirillov et al., 2023). Related efforts such as Change-UP (Yang et al., 2025b) and ChangeMinds (Wang et al., 2024) target joint RSCDC within unified architectures, end-to-end architectures. Change-UP emphasises multi-scale change-aware feature fusion and semantic query alignment to enhance differential representations. ChangeMinds, by contrast, introduces a change-aware LSTM and a cross-attention predictor to learn shared spatiotemporal features for both change detection and captioning. However, these approaches - whether task-specific or unified - remain narrow in objective scope, relying on specialised labelled datasets that limit generalisation (e.g. RSGPT (Hu et al., 2025)) (Liu et al., 2025). Another direction aims to unify RS tasks within broader vision-language foundation models, moving beyond task-specific adaptations to capture general visual-textual representations of Earth observation data across space and time (Li et al., 2024d). GeoChat (Kuckreja et al., 2024) and EarthGPT (Zhang et al., 2024b) adapt LLaVa-style conversational VLMs for RS multi-task reasoning, with GeoChat supporting grounded region-aware dialogue and EarthGPT extending to multi-sensor interpretation and unified RS task handling. RSUniVLM (Liu and Lian, 2024) advances this approach by unifying image-, region-, and pixel-level understanding through a granularity-oriented mixture-of-experts architecture, enabling both fine-grained spatial reasoning and multi-image tasks. UniRS (Li et al., 2024d) further addresses the single-image limitation by introducing a unified spatiotemporal visual representation and a dedicated change-extraction module, allowing a single model to process single images, bi-temporal pairs, and video sequences for comprehensive multi-temporal RS analysis.
@@
Most recently, LLM-driven RS vision–language agents extend these models by integrating perception, reasoning, and tool use into multistep analytical workflows, supporting tasks such as temporal scene understanding and change detection through multi-turn conversations (Irvin et al., 2025; Liu et al., 2024). Tree-GPT is an example of an LLM-driven RS vision-language agent for forest scene analysis. It combines LLM orchestration with knowledge bases, vision understanding modules and a web UI to support interactive forest scene analysis, albeit lacking a temporal component.
@@ 2.4 LLM Agents for Remote Sensing
AI agents are systems that link perception and action through iterative reasoning to achieve user-defined goals (Xu et al., 2024). Utilising LLMs pretrained on massive text corpora, they can learn from interactions, decompose tasks, and interact dynamically across steps with external data, models, or tools (Liu et al., 2025), serving as suitable modular orchestrators for complex workflow execution (Liu et al., 2024; Schick et al., 2023; Yao et al., 2023; Guo et al., 2024).
@@
Within RS, LLM-based agents are still in their infancy, with applications for Earth observation tasks gradually increasing (Xu et al., 2024; Liu et al., 2024; Du et al., 2023). They are typically task-specific and rely on strict prompt templates and few-shot examples to guide reasoning and improve robustness (Xu et al., 2024; Guo et al., 2024). Agents execute tasks by selecting toolkits that often contain specialised visual modules, which can be run independently or composed into sequential pipelines to accomplish multiple objectives (Guo et al., 2024; Du et al., 2023), as in Change-Agent (Liu et al., 2024) and RS-Agent (Xu et al., 2024). This modular design, which often benefits from integrated knowledge bases and elucidation of task functionalities, helps reduce hallucinations, enhances accuracy, and enables flexible, customisable outputs (Guo et al., 2024; Yao et al., 2023; Liu et al., 2025; Du et al., 2023). RS-specific LLM agents generally fall into two broad paradigms. The first comprises modular tool-using systems that rely on external LLMs to orchestrate specialised vision modules. The second consists of end-to-end instruction-tuned VLMs that embed temporal or multimodal reasoning directly within the model. Recent work illustrates both directions (Liu et al., 2024; Xu et al., 2024; Deng et al., 2025), including hybrid approaches (Irvin et al., 2025).
@@
When introducing ChangeChat - an interactive VLM designed for bi-temporal RS change analysis using multimodal instruction tuning - Deng et al. (2025) claim that approaches such as Change-Agent (Liu et al., 2024) and RS-ChatGPT (Guo et al., 2024), that rely on external LLMs, are limited in end-to-end flexibility, with system capabilities constrained to pre-defined modules. TEOChat (Irvin et al., 2025) bridges the two design philosophies by incorporating instruction tuning over variable length temporal image sequences, enabling stepwise reasoning across multi-date satellite imagery while maintaining modular interpretability. Another prominent limitation raised by Xu et al. (2024) is the challenge of providing flexible task orchestration and mitigating failures in task decomposition reasoning and tool selection, as witnessed with models like RS-Chatgpt (Guo et al., 2024). RS-Agent addresses this by using an LLM as a controller, which parses user instructions, selects from a suite of geospatial tools, and composes multistep workflows. A retrieval-augmented knowledge graph supports the LLM with domain-specific information and task-aware strategies to enable complex reasoning across diverse RS tasks. Change-Agent (Liu et al., 2024), is designed specifically for bi-temporal change analysis, integrating an LLM with change-focused reasoning modules to support interactive and interpretable assessments of temporal differences in RS imagery.
@@
Comparing the capabilities of these various LLM-based RS agents is challenging due to differences in task scope, input modalities, architectural configuration, and evaluation setups. This lack of standardised benchmarking makes it difficult to assess performance, task orchestration, and agentic reasoning across approaches. To date, only ChangeMinds (Wang et al., 2024) and Change-UP (Yang et al., 2025b) benchmark against Change-Agent’s MCI model (Liu et al., 2024) for joint RSCD and RSICC, with both reporting modest gains. However, neither work incorporates an LLM-based agent, or releases code. Other framework either lack temporal reasoning (Guo et al., 2024; Du et al., 2023; Xu et al., 2024), or are limited in their CDC capabilities (Irvin et al., 2025; Deng et al., 2025), leaving Change-Agent as the most appropriate framework identified.
@@ 3.1 LEVIR-MCI-Trees
When developing the robust MCI model at the heart of the Change-Agent system, Liu et al. (2024) proposed the LEVIR-MCI dataset, which comprises both pixel-level change information contained within change detection masks and descriptive semantic-level captions. Each of the 10,077 examples contains bi-temporal image pairs with a variable time span of 51̃5 years, a corresponding annotated mask, and five annotated captions. Each image spans $\times256$ pixels at a high spatial resolution of 0.5m/pixel. LEVIR-MCI extends LEVIR-CC (Liu et al., 2022) by providing each pair of bi-temporal images with change detection masks that explicitly highlight changes to roads and buildings, with over 40,000 changed roads and buildings annotated. The five captions for each image pair are intended to provide diverse annotations from varying perspectives to bolster its utility for change detection.
For the purposes of this study, a subset of this dataset was taken that discards image pairs that contain no changes to tree cover, hereby referred to as LEVIR-MCI-Trees. Filtering is performed based on the contents of the captions for each example. If none of the captions for an example contains one of: ‘tree, trees, wood, woods, woodland, wooded, forest, forests, jungle, jungles‘ then it is removed. Post filtering, the resulting subset contains 2,305 examples distributed across training, validation and test sets comprising 66%, 16% and 18% (1518, 374, 413) of the dataset respectively. Due to the dataset only containing pixel-level annotations for roads and buildings, it is unsuitable for benchmarking Forest-Chat’s performance for deforestation segmentation directly, but can be used to gauge the architecture’s generalisation capabilities to new geographical domains. However, this subset will be suitable for comparing the performance of tree disturbance captioning and highlighting any differences in the difficulty of pixel-level annotation between urban and forest contexts. Segmentation performance in urban settings is generally higher than in natural landscapes due to simpler object geometries and clearer boundaries (Wang et al., 2025), leading to a preference in benchmarking model performance within urban contexts.
@@ 3.2.1 Data Curation
Figure 2: Summary statistics of change cover in segmentation masks for Forest-Change and LEVIR-MCI-Trees datasets.
@@ 3.2.2 Preprocessing
The Forest-Change dataset is organised to mirror the directory structure of LEVIR-MCI-Trees to maintain consistent data loading. Three splits were creating for training, validation, and testing at an approximate ratio of 80%:10%:10% (270, 31, and 33 samples, respectively). All images are resized to $\times256$ pixels to match LEVIR-MCI-Trees, and pixel-wise change masks are binarised to indicate change (1) or no-change (0). The bi-temporal image pairs are pre-aligned, and no additional geometric registration or atmospheric correction was applied. Images were normalised using dataset-specific per-channel mean and standard deviation statistics. No additional radiometric correction or cloud masking was performed, although a few samples exhibit partial cloud occlusion.
@@ 3.2.3 Caption Generation
As shown in Figure 2, captions for the Forest-Change dataset are skewed toward descriptions of lower degrees of forest loss and recurring loss patterns. This reflects the limited scene diversity introduced by the cropping-augmentation strategy adopted by Hewarathna et al. (2024), as discussed previously in Section 3.2.1. Captioning extensive amounts of remote sensing change imagery is fatiguing for human annotators, especially when underlying scene diversity is limited. Therefore, rule-based caption generation is proposed as a way to ensure minimal requirements for adequate semantic context and sentence structure are met. Human annotations are still valuable for improving linguistic and semantic variation, and providing geographic context.
@@
For automated caption generation, the percentage of newly deforested area, the size of individual change patches, and their spatial distribution are computed from the mask. The percentage of forest loss is binned into adjective-based descriptors aligned with the model vocabulary. A rule-based caption generator was designed to introduce lexical and syntactic variation, producing complementary captions to the human-provided description. Although LLMs are increasingly used for synthetic text generation in NLP tasks Yu et al. (2023), they were not adopted here due to the risk of introducing prompt-induced bias and the limited scope of the project in which to mitigate it.
@@ 4 Methodology
This section outlines the components of the Forest-Chat agent. The goal of the Forest-Chat agent is to provide a conversational interface for bi-temporal remote sensing (RS) forest change analysis through joint pixel-level change detection and change captioning (CDC). Building on the Change-Agent framework (Liu et al., 2024), Forest-Chat integrates a domain-specific VLM (the MCI model) for supervised CDC, a zero-shot foundation vision model (AnyChange) for class-agnostic pixel-level change localisation, and an LLM controller for instruction interpretation, tool selection, and analytical reasoning. To adapt Forest-Chat for forest change analysis, the MCI model is trained on the Forest-Change and LEVIR-MCI-Trees datasets to produce structured change masks and semantic change descriptions. When used in a supervised capacity, the system is referred to as FC-Supervised. The system is further augmented with the AnyChange model to enable zero-shot change detection on out-of-distribution data, with this capability termed as FC-Zero-shot. To support interactive zero-shot change analysis, an interface is developed for user-provided point prompts to guide object-oriented change detection with AnyChange. Finally, the Forest-Chat toolset is extended with forest-specific analytical functions, including deforestation area estimation and patch-level statistics, enabling both low-level perception and high-level reasoning within a unified agentic framework.
The overall system diagram for Forest-Chat is presented in Figure 1, while Figure 5 provides a simplified overview of the perception models incorporated into Forest-Chat. The components of Forest-Chat are described below.
@@
Figure 5: Overview of the two visual perception models available to the Forest-Chat agent. (a) highlights the overall structure of the MCI model (b) introduces a simplified overview of the AnyChange model.

### v2

@@ 2.2 Vision-Language Models for Remote Sensing
Vision–language models (VLMs) combine instruction-following with visual feature analysis, typically pairing a pretrained visual encoder - such as CLIP model (Radford et al., 2021), or a ViT - with an LLM (e.g. GPT (openai et al., 2023), LLaMa (Touvron et al., 2023)) through a projection or transformation module that maps visual embeddings into the language model’s input space (Li et al., 2024d; Tao et al., 2025). Recent work has explored improved vision-language pretraining strategies for aligning visual and textual representations (Qin et al., 2025). This alignment enables generalised tasks such as visual question answering, captioning, and zero-shot transfer learning (Lu et al., 2025; Li et al., 2024b). These capabilities have made VLMs an increasingly common choice for RS applications (Zhang et al., 2025; Liu and Lian, 2024; Li et al., 2024d; Hu et al., 2025; Zhang et al., 2024b), which benefit from domain-specific pretraining across task-agnostic objectives, varying sensors, resolutions, and imaging geometries, including multimodal fusion of complementary data sources (Yang et al., 2024; Lu et al., 2025). Self-supervised and zero-shot learning further reduces reliance on exhaustive annotation (Peng et al., 2025; Lu et al., 2025), while text prompting and question-answering interactions support context-aware interpretation, helping models link ambiguous visual patterns to high-level semantic concepts such as deforestation (Li et al., 2024a).
@@
Despite strong performance on natural imagery, general-purpose multimodal LLMs (MLLMs) such as GPT-4V and LLaVa (openai et al., 2023; Bazi et al., 2024) underperform on several RS and change-related tasks due to domain mismatch (Zhang and Wang, 2024). One response has been to adapt LLM-driven, task-specific spatio-temporal VLMs to RS vision-language understanding tasks (Liu et al., 2025a), including models such as KCFI (Yang et al., 2025a) and Semantic-CC (Zhu et al., 2024), that fine-tune visual modules or lightweight LLM components for change-focused tasks. Related efforts such as Change-UP (Yang et al., 2025b) and ChangeMinds (Wang et al., 2024) target joint RSCDC within unified, end-to-end architectures. However, these approaches remain narrow in objective scope, relying on specialised labelled datasets that limit generalisation (e.g. RSGPT (Hu et al., 2025)) (Liu et al., 2025a). A broader direction aims to unify RS tasks within vision-language foundation models (Li et al., 2024d). GeoChat (Kuckreja et al., 2024) and EarthGPT (Zhang et al., 2024b) adapt LLaVa-style conversational VLMs for multi-task, grounded region-aware dialogue and multi-sensor interpretation while RSUniVLM (Liu and Lian, 2024) and UniRS (Li et al., 2024d) extend this to pixel-level understanding and spatio-temporal multi-image analysis respectively.
@@
Most recently, LLM-driven RS vision–language agents extend these models by integrating perception, reasoning, and tool use into multistep analytical workflows, supporting tasks such as temporal scene understanding and change detection through multi-turn conversations (Irvin et al., 2025; Liu et al., 2024). Tree-GPT (Du et al., 2023) is one such example, combining LLM orchestration with knowledge bases and vision modules for interactive forest scene analysis, albeit lacking a temporal component.
@@ 2.3 LLM Agents for Remote Sensing
AI agents link perception and action through iterative reasoning to achieve user-defined goals (Xu et al., 2024). Utilising LLMs pretrained on massive text corpora, they can decompose tasks and interact dynamically across steps with external data, models, or tools (Liu et al., 2025a), serving as modular orchestrators for complex workflow execution (Liu et al., 2024; Schick et al., 2023; Yao et al., 2023; Guo et al., 2024).
@@
Within RS, LLM-based agents are still in their infancy (Xu et al., 2024). They are typically task-specific, relying on strict prompt templates and few-shot examples to guide reasoning (Guo et al., 2024). Agents execute tasks by selecting from specialised visual toolkits, composable into sequential pipelines, as in Change-Agent (Liu et al., 2024) and RS-Agent (Xu et al., 2024) to accomplish multiple objectives. This modular design - often augmented with knowledge bases - reduces hallucinations and enables flexible outputs (Guo et al., 2024; Liu et al., 2025a). RS-specific LLM agents generally fall into two broad paradigms: modular tool-using systems that rely on external LLMs to orchestrate specialised vision modules; and end-to-end instruction-tuned VLMs that embed temporal or multimodal reasoning directly within the model (Liu et al., 2024; Xu et al., 2024; Deng et al., 2025; Irvin et al., 2025).
@@
Key limitations span both paradigms - modular systems face restricted end-to-end flexibility (Deng et al., 2025), while flexible task orchestration, mitigating task decomposition failures, and robust tool selection remain challenging across approaches (Xu et al., 2024). TEOChat (Irvin et al., 2025) bridges these design philosophies through instruction tuning over variable-length temporal image sequences, enabling stepwise reasoning across multi-date imagery while maintaining modular interpretability. RS-Agent (Xu et al., 2024) addresses these issues by using an LLM controller that selects from a suite of geospatial tools and composes multistep workflows, supported by a retrieval-augmented knowledge graph for domain-specific reasoning. Change-Agent (Liu et al., 2024) is designed specifically for bi-temporal change analysis, integrating an LLM with change-focused reasoning modules to support interactive and interpretable assessments of temporal differences in RS imagery.
@@
Comparing these agents is challenging due to differences in task scope, input modalities, and evaluation setups, making it difficult to assess performance and agentic reasoning across approaches (Shabbir et al., 2025). To date, only ChangeMinds (Wang et al., 2024) and Change-UP (Yang et al., 2025b) benchmark against Change-Agent’s MCI model (Liu et al., 2024) for joint RSCD and RSICC, with both reporting modest gains. However, neither incorporates an LLM-based agent, or is publicly available. Other frameworks either lack temporal reasoning (Guo et al., 2024; Du et al., 2023), or offer limited CDC capabilities (Irvin et al., 2025; Deng et al., 2025), leaving Change-Agent as the most appropriate baseline identified.
@@ 3.1 LEVIR-MCI-Trees
When developing the robust MCI model at the heart of the Change-Agent system, Liu et al. (2024) proposed the LEVIR-MCI dataset, which comprises both pixel-level change information contained within change detection masks and descriptive semantic-level captions. Each of the 10,077 examples contains bi-temporal image pairs with a variable time span of 5-15 years, a corresponding annotated mask, and five annotated captions. Each image spans $\times256$ pixels at a high spatial resolution of 0.5m/pixel. LEVIR-MCI extends LEVIR-CC (Liu et al., 2022a) by providing each pair of bi-temporal images with change detection masks that explicitly highlight changes to roads and buildings, with over 40,000 changed roads and buildings annotated. The five captions for each image pair are intended to provide diverse annotations from varying perspectives to bolster its utility for change detection.
For the purposes of this study, a subset of this dataset was taken that discards image pairs that contain no changes to tree cover, hereby referred to as LEVIR-MCI-Trees. Filtering is performed based on the contents of the captions for each example. If none of the captions for an example contains one of: ‘tree, trees, wood, woods, woodland, wooded, forest, forests, jungle, jungles‘ then it is removed. Post filtering, the resulting subset contains 2,305 examples distributed across training, validation and test sets comprising 66%, 16%, and 18% (1518, 374, 413) of the dataset respectively. Due to the dataset only containing pixel-level annotations for roads and buildings, it is unsuitable for benchmarking Forest-Chat’s performance for deforestation segmentation directly, but can be used to gauge the architecture’s generalisation capabilities to new geographical domains. However, this subset will be suitable for comparing the performance of tree disturbance captioning and highlighting any differences in the difficulty of pixel-level annotation between urban and forest contexts. Segmentation performance in urban settings is generally higher than in natural landscapes due to simpler object geometries and clearer boundaries (Wang et al., 2025), leading to a preference in benchmarking model performance within urban contexts.
@@ 3.2.1 Data Curation
Figure 2: Summary statistics of change cover in segmentation masks for the Forest-Change, LEVIR-MCI-Trees, and JL1-CD-Trees datasets.
@@ 3.2.2 Preprocessing
The Forest-Change dataset is organised to mirror the directory structure of LEVIR-MCI-Trees to maintain consistent data loading. Three splits were created for training, validation, and testing at an approximate ratio of 80%:10%:10% (270, 31, and 33 samples, respectively). All images are resized to $\times256$ pixels to match LEVIR-MCI-Trees, and pixel-wise change masks are binarised to indicate change (1) or no-change (0). The bi-temporal image pairs are pre-aligned, and no additional geometric registration or atmospheric correction was applied. Images were normalised using dataset-specific per-channel mean and standard deviation statistics. No additional radiometric correction or cloud masking was performed, although a few samples exhibit partial cloud occlusion.
@@ 3.2.3 Caption Generation
As shown in Figure 3, captions for the Forest-Change dataset are skewed toward descriptions of minor forest loss and recurring loss patterns. This reflects the limited scene diversity introduced by the cropping-augmentation strategy adopted by Hewarathna et al. (2024), as discussed previously in Section 3.2.1. Captioning large volumes of remote sensing change imagery is labour-intensive for human annotators, especially when scene diversity is limited. Therefore, rule-based caption generation is proposed as a way to ensure minimal requirements for semantic context and sentence composition are met. Human annotations remain valuable for improving linguistic and semantic variation, and providing geographic context.
@@
For automated caption generation, the percentage of newly deforested area, the size of individual change patches, and their spatial distribution are computed from the mask. The percentage of forest loss is binned into adjective-based descriptors aligned with the model vocabulary. A rule-based caption generator was designed to reduce labelling burden, and enforce sufficient change descriptions are generated, producing complementary captions to the human-provided ones. Although LLMs are increasingly used for synthetic text generation in NLP tasks Yu et al. (2023), they were not adopted here due to the risk of introducing prompt-induced bias and the limited scope of the project in which to mitigate it.
@@
3.3 JL1-CD-Trees
Remote sensing datasets often contain varying degrees of noise, seasonal variability, cross-sensor inconsistencies, and domain shifts. To evaluate Forest-Chat’s robustness and domain transfer capabilities for unseen forest change detection scenarios, we introduce JL1-CD-Trees. The JL1-CD dataset was proposed by Liu et al. (2025b) as a benchmark for evaluating CD algorithms. It comprises 5,000 satellite image pairs captured by the Jilin-1 satellite, across various regions of China between early 2022 and the end of 2023. The imagery has sub-meter spatial resolution ranging from 0.5 to 0.75 meters. The dataset includes a diverse set of anthropogenic and natural surface features, including buildings, roads, hardened surfaces, woodlands, grasslands, croplands, water bodies, and photovoltaic (PV) panels. Each change mask may contain multiple surface change types.
To create JL1-CD-Trees, the dataset was manually filtered to only retain samples containing changes between woodland and other natural surface types (e.g. grassland), thereby maximising domain alignment with Forest-Change. The resulting subset presents several challenges, including seasonal variation, atmospheric artifacts (e.g. cloud cover and haze), and both gains and losses in tree cover. After filtering, the dataset contains 408 samples, split into training, validation, and test sets with an approximate 60:20:20 ratio (244, 81, and 83 samples, respectively). To ensure consistency with Forest-Change, identical preprocessing steps are applied, including image rescaling, change mask binarisation, and pixel value normalisation. Figure 2 highlights that JL1-CD-Trees and Forest-Change have similar data distributions and class imbalances. JL1-CD-Trees is intended solely for evaluating domain transfer in forest change detection; therefore no captions are provided.
@@ 4 Methodology
This section outlines the components of the Forest-Chat agent. The goal of the Forest-Chat agent is to provide a conversational interface for bi-temporal RS forest change analysis through joint pixel-level change detection and change captioning (CDC). Building on the Change-Agent framework (Liu et al., 2024), Forest-Chat integrates a domain-specific VLM (the MCI model) for supervised CDC, a zero-shot foundation model (AnyChange) for class-agnostic pixel-level change localisation, and GPT-4o for zero-shot change captioning and refinement. These supervised and zero-shot configurations are referred to as FC-Supervised and FC-Zero-shot respectively. To adapt Forest-Chat for forest change analysis, the MCI model is trained on the Forest-Change and LEVIR-MCI-Trees datasets to produce structured change masks and semantic change descriptions. An interactive interface is additionally developed for user-provided point prompts to guide object-oriented change detection with AnyChange. An LLM controller oversees instruction interpretation, tool selection, and analytical reasoning across all modules. The system is further augmented with forest-specific analytical functions, including deforestation area estimation and patch-level statistics, enabling both low-level perception and high-level reasoning within a unified agentic framework.
The overall system diagram is presented in Figure 1, with Figure 5 providing a simplified overview of the incorporated perception models.
@@
Figure 5: Overview of the two visual perception models available to the Forest-Chat agent. (a) Highlights the overall structure of the MCI model (b) Introduces a simplified overview of the AnyChange model. GPT-4o is additionally integrated for zero-shot captioning and captioning refinement, but a diagram is not provided due to its proprietary status.